ahhh help - I lost my notes. I got (17,5) but how?
4x-y=3
9x-2y=11

Answer by stanbon(75887)

4x-y=3
9x-2y=11
-------
Multiply thru the top equation by 2:
8x -2y = 6
9x -2y = 11
--------------
Subtract and solve for "x":
x = 5
------
Solve for "y":
4x - y = 3
4*5 - y = 3
y = 17
====================
Cheers,
Stan H.